# NExON-Bayes: A Bayesian approach to network estimation informed by ordinal covariates

### v1

@@ Abstract
Results: Our modelling framework, called NExON-Bayes, extends the graphical spike-and-slab framework to account for ordinal covariates, jointly estimating their relevance to the graph structure and leveraging them to improve the accuracy of network estimation. To scale to high-dimensional omic settings, we develop an efficient variational inference algorithm tailored to our model. Through simulations, we demonstrate that our method outperforms the vanilla graphical spike-and-slab (with no covariate information), as well as other state-of-the-art network approaches which exploit covariate information. Applying our method to reverse phase protein array data from patients diagnosed with stage I, II or III breast carcinoma, we estimate the behaviour of proteomic networks as breast carcinoma progresses. Our model provides insights not only through inspection of the estimated proteomic networks, but also of the estimated ordinal covariate dependencies of key groups of proteins within those networks, offering a comprehensive understanding of how biological pathways shift across disease stages.
@@ 1 Introduction
More recently, methods have been developed to jointly estimate graphical models across multiple networks, reflecting such conditions. Joint methods estimate network structures through sharing information between networks when appropriate, while ideally retaining network-specific differences. Several approaches for joint estimation of multiple GGMs have been proposed. Frequentist approaches extend the penalised likelihood framework to add penalty parameters that encourage similarity between network-specific precision matrices (Guo et al., 2011; Ma and Michailidis, 2016; Lingjærde and Richardson, 2023), whilst Bayesian approaches use priors to encourage the potential sharing of network structures (Li et al., 2019; Lingjærde et al., 2024). Such methods have been used, for example, to estimate the proteomic networks of breast cancer patients with different subtype diagnoses, where the expected similarities between networks are leveraged in the joint estimation (Lingjærde and Richardson, 2023).
@@
Existing joint estimation approaches often treat networks as exchangeable, which in some cases can inhibit statistical power. For example, when the covariates used to define subgroups have a natural ordering, the assumption of exchangeability both omits useful information that could be leveraged in the estimation and limits the potential insight into how those covariates affect graphical structure. A related set of more recent approaches build on this idea by making GGMs dependent on covariates. Despite working for a wide range of covariates, these approaches do not ensure the positive definiteness of the precision matrix and often lack scalability, which seems to hamper widespread adoption (Helwig et al., 2024; Chen et al., 2024).
@@
Here we present NExON-Bayes, a Bayesian approach to network estimation informed by ordinal covariates. NExON-Bayes builds on the standard graphical spike-and-slab model of Wang et al. (2015) Wang (2015), and introduces a sub-model that characterises the dependence between edge inclusion probabilities and sample-level ordinal covariate data, rather than modelling the entries of the precision matrix, as done in other covariate-dependent approaches. This modelling framework allows us to develop a deterministic inference algorithm that is more scalable than existing alternatives in the literature.
@@
This paper is organised as follows. Section 2 introduces the terminology and notation used throughout the article, our proposed model, and the variational inference algorithm. Next, Section 3 assesses the performance of our approach by benchmarking the model against several competitors through simulations, and investigates its benefits in a real-life setting where we estimate the proteomic networks of breast carcinoma patients. Finally, Section 4 summarises our contributions and avenues for extensions of this work.
@@ 2.1 Terminology and notation
where $\boldsymbol{\Omega}$ is the positive definite inverse covariance matrix, known as the precision matrix. In this work, we consider the case in which each of the N samples is characterised by a sample-level ordinal covariate $a$ taking values in a set $\mathcal{A}$. We further assume sample homogeneity within groups defined by the same covariate value, rather than across all samples. Specifically,
@@
where $\boldsymbol{\Omega}^{(a)}$ is the precision matrix specific to group $a\in\mathcal{A}$, [^1] When referring to the covariate value, we use ‘$a$’, and when referring to networks or data that corresponds to a covariate value $a$ we use ‘$(a)$’. and $N_{a}$ is the number of samples within the dataset specific to a covariate value $a$. A covariate-specific dataset can then be denoted as $\boldsymbol{Y}^{(a)}=(\boldsymbol{y}^{(a)}_{1},...,\boldsymbol{y}^{(a)}_{N_{a}})$.
@@ 2.2 Graphical spike-and-slab model
To induce sparsity, our proposed approach, NExON-Bayes, relies on a continuous spike-and-slab prior on the off-diagonal elements of each $\mathbf{\Omega}^{(a)}$, similarly as in Wang et al. (2015) Wang (2015). This is achieved by using a mixture of two Gaussian distributions, one with very low variance (the spike) and the other with high variance (the slab). A binary edge inclusion indicator is also introduced within this prior, which is denoted as $\delta_{ij}^{(a)}$, corresponding to the $ij$’th entry of $\mathbf{\Omega}^{(a)}$. Edge inclusion is indicated by $\delta_{ij}^{(a)}=1$, while edge exclusion is indicated by $\delta_{ij}^{(a)}=0$. The spike-and-slab prior is thus represented as
@@
for $1\leq i<j\leq P$, $i\neq j$. A symmetric matrix is formed by mirroring this lower triangle across the diagonal.
@@ 2.3 Dependence on ordinal covariates
where $\Phi(\cdot)$ is the standard normal cumulative distribution function. The parameter $\beta_{ij}$ is an edge-specific regression coefficient quantifying the influence of the covariate on the inclusion of edge $(i,j)$. Specifically, $\beta_{ij}=0$ means no relationship between covariate value and edge inclusion probability, i.e., edge $(i,j)$ has the same probability of being present or absent in all networks. A positive value of $\beta_{ij}$ means that edge inclusion is more likely in networks with high covariate values and vice versa. In addition, $\zeta_{ij}$ captures the varying baseline inclusion probability of edges and acts as an offset within the probit submodel. We assign a Gaussian prior distribution to $\beta_{ij}$,
@@
Each $\zeta_{ij}$ is assigned a Gaussian prior,
@@
where the parameters $n_{0}$ and $t_{0}^{2}$ are fixed following Xi and Ruffieux (2024) to prior beliefs about the mean and variance of the number of edges in the networks, and thus the overall sparsity.
@@ 2.4 Variational inference algorithm
Stochastic sampling algorithms like Markov chain Monte Carlo (MCMC) are typically computationally demanding for large graphical models (Wang, 2015). Instead, we employ deterministic inference algorithms which have been successfully applied in graphical modelling contexts, see, e.g., Li and McCormick (2019), Lingjærde et al. (2024), Xi and Ruffieux (2024). Specifically, we develop a variational Bayes expectation conditional maximisation (VBECM) algorithm. The true posterior distribution can be written as $p(\underline{\boldsymbol{\Omega}},\boldsymbol{\Theta}\mid\boldsymbol{y})$, where $\underline{\boldsymbol{\Omega}}=\{\boldsymbol{\Omega}^{(a)}\}_{a\in\mathcal{A}}$ and $\boldsymbol{\Theta}$ is the set of all variables outside of the precision matrix. This true distribution is approximated by using a variational distribution with mean-field approximations:
$q(\underline{\boldsymbol{\Omega}},\boldsymbol{\Theta})$ (2)
$=\prod_{a\in\mathcal{A}}q(\boldsymbol{\Omega}^{(a)})\prod_{a\in\mathcal{A}}\prod_{i<j}q(\delta^{(a)}_{ij},z_{ij}^{(a)})\prod_{i<j}q(\zeta_{ij})q(\sigma^{-2})\prod_{i<j}q(\beta_{ij}),$
@@
$\mathcal{L}(q)=\mathbb{E}_{q(\underline{\boldsymbol{\Omega}},\boldsymbol{\Theta})}\{\log p(\boldsymbol{Y},\underline{\boldsymbol{\Omega}},\boldsymbol{\Theta})\}-\mathbb{E}_{q(\underline{\boldsymbol{\Omega}},\boldsymbol{\Theta})}\{\log q(\underline{\boldsymbol{\Omega}},\boldsymbol{\Theta})\},$
(Bishop and Nasrabadi, 2006). The optimal distribution for the $n$-th factor of $\boldsymbol{\Theta}$, denoted $q_{n}(\boldsymbol{\Theta}_{n})$ is found by iteratively taking the expectation of the logarithm of the joint distribution over all the other factors,
@@
where the constant absorbs all terms not depending on $\boldsymbol{\Theta}_{n}$. However, the variational distribution of each $\boldsymbol{\Omega}^{(a)}$ is intractable; we therefore update $\underline{\boldsymbol{\Omega}}$ blockwise using a conditional-maximisation step according to Wang (2015).
Finally, we use the extended Bayesian Information Criteria ($\text{BIC}_{\gamma}$) to select the spike variance $\nu_{0}$. $\text{BIC}_{\gamma}$ extends the Bayesian Information Criteria (BIC) with an additional regularisation penalty parameter $\gamma\in[0,1]$. The $\text{BIC}_{\gamma}$ of a network-specific graphical estimation, $\boldsymbol{\Omega}^{(a)}_{est.}$, can be found using the following equation:
@@
where $\ell(\cdot)$ is the log likelihood and $|E^{(a)}|$ is the number of edges present in estimated network $(a)$ (Foygel and Drton, 2010). The regularisation parameter $\gamma$ can be tuned to control the sparsity of the estimates. In this paper, we use the default value of $\gamma=0.5$.
To optimise the values of $\nu_{0}$, we use line searches to find a value of $\nu_{0}^{(a)}$ that minimises $\text{BIC}_{\gamma}(\boldsymbol{\Omega}_{est.}^{{(a)}})$ for each network $(a)$ in the vanilla graphical spike-and-slab model, which gives a strong estimate for expected sparsity.
@@ 3.1 Simulated Data
We use simulation studies to assess the performance of NExON-Bayes by analyzing how well it can recover the true precision matrix from which simulated data is generated. For this setting, we consider an ordinal covariate taking values $a\in\mathcal{A}=\{1,2,3,4\}$. We correspondingly generate $|\mathcal{A}|=4$ precision matrices, each with $P=100$ nodes, from which we sample normally distributed data according to Equation (1), where the number of observations for each dataset is $N=150$. To start, network $(1)$ is simulated as scale-free, with $100$ edges ($|E|=P$) and thus a sparsity of $0.02$. ‘Scale-free’ refers to a graph that has a degree distribution that follows a power law, which leads to few nodes having several connected edges (hubs) and most nodes having very few connected edges. This trait emulates generic biological networks, which are often scale-free (Barabási and Bonabeau, 2003). The other networks are then created by altering the entries of this first network. Specifically, the zero-valued partial correlations corresponding to $50$ edges that are not present in this network $(1)$ are simulated to linearly increase with the value of the covariate $a$. Similarly, the non-zero valued partial correlations corresponding to $50$ edges that are present in this network $(1)$ are simulated to linearly diminish to zero, such that they are not present in network $(4)$. The remaining $50$ edges present in network $(1)$ are present in all four networks, with unchanged partial correlation values. We call edges with increasing partial correlation “appearing” edges and those with decreasing partial correlation “disappearing” edges.
Simulating precision matrices in this way gives rise to an interesting artefact: appearing and disappearing edges have smaller precision matrix entries in networks $(2)$ and $(3)$ and are thus more difficult to uncover. These networks are shown schematically in Figure S1 and Section 2.1 of the Supplementary Material which also provides further details of the data generation algorithm.
@@ 3.2 Borrowing information across networks using joint modelling
Figure 1: Posterior mean estimates of each $\beta_{ij}$ (lower diagonal part) versus simulated coefficients of the ordinal covariate (upper diagonal part). Note that there is no real “true" $\mathbf{\beta}$, as the precision matrices is not simulated from the model. However, there is a known structure based on positive ($\beta_{ij}>0$) and negative ($\beta_{ij}<0$) dependencies between covariate and partial correlation. The outline of the true structure is overlain on the estimated structure to ensure a clear comparison (black).
We now use this simulated data to compare the performance of NExON-Bayes to the standard single-network Bayesian graphical spike-and-slab model (SSL) (Xi and Ruffieux, 2024; Li and McCormick, 2019), upon which our extension is built. Specifically, we evaluate whether introducing a direct relationship between a covariate and the posterior probability of edge inclusion (PPI) indeed boosts statistical power – as a first sanity check that the model behaves as it is meant to.
@@
As discussed earlier, the level of information sharing is controlled by the edge-specific regression coefficient $\beta_{ij}$. Not only is the estimation of $\beta_{ij}$ useful in increasing statistical power, it also provides insight on sample-level heterogeneity in terms of the effect of the sample-level covariate on the graphical structure. We now consider a smaller simulated setting (with $P=20$) focused on assessing the ability of NExON-Bayes to recover the values of $\beta_{ij}$. The results, shown in Figure 1, indicate that NExON-Bayes successfully captures the changes of edges associated with covariates, information which, as we will illustrate in the cancer data analysis, can enable clinically meaningful interpretation of the estimated network structure. Such insights are not captured intrinsically by state-of-the-art single or joint network modelling approaches.

### v2

@@ Abstract
Results: Our modelling framework, called NExON-Bayes, extends the graphical spike-and-slab framework to account for ordinal covariates, jointly estimating their relevance to the graph structure and leveraging them to improve the accuracy of network estimation. To scale to high-dimensional omic settings, we develop an efficient variational inference algorithm tailored to our model. Through simulations, we demonstrate that our method outperforms the vanilla graphical spike-and-slab (with no covariate information), as well as other state-of-the-art network approaches which exploit covariate information. Applying our method to reverse phase protein array data from patients diagnosed with stage I, II or III breast carcinoma, we estimate the behaviour of proteomic networks as cancer progresses. Our model provides insights not only through inspection of the estimated proteomic networks, but also of the estimated ordinal covariate dependencies of key groups of proteins within those networks, offering a comprehensive understanding of how biological pathways shift across disease stages.
@@ 1 Introduction
More recently, methods have been developed to jointly estimate graphical models across multiple networks, reflecting such conditions. Joint methods estimate network structures through sharing information between networks when appropriate, while ideally retaining network-specific differences. Several approaches for joint estimation of multiple GGMs have been proposed. Frequentist approaches extend the penalised likelihood framework to add penalty parameters that encourage similarity between network-specific precision matrices (Guo et al., 2011; Ma and Michailidis, 2016; Lingjærde and Richardson, 2023), whilst Bayesian approaches use priors to encourage the potential sharing of network structures [Peterson et al., 2015, Li et al., 2019; Lingjærde et al., 2024]. Such methods have been used, for example, to estimate the proteomic networks of breast cancer patients with different subtype diagnoses, where the expected similarities between networks are leveraged in the joint estimation (Lingjærde and Richardson, 2023).
@@
Existing joint estimation approaches often treat networks as exchangeable, which in some cases can inhibit statistical power. For example, when the covariates used to define subgroups have a natural ordering, the assumption of exchangeability both omits useful information that could be leveraged in the estimation and limits the potential insight into how those covariates affect graphical structure. A related set of more recent approaches build on this idea by making GGMs dependent on covariates. Despite being applicable in a plethora of settings, these approaches do not ensure the positive definiteness of the precision matrix and often lack scalability, which seems to hamper widespread adoption (Ni et al., 2022; Niu et al., 2024).
@@
Here we present NExON-Bayes, a Bayesian approach to network estimation informed by ordinal covariates. NExON-Bayes is thus applicable to settings with explicit ordinal covariates or where an ordinal structure can reasonably assumed. These covariates encode auxiliary, non-omic data, which can represent a varied information, such as tumour progression with a covariate taking values in $\{1,2,3\}$ corresponding to stage I, II & III cancers, therapeutic response categories with covariate taking values in $\{1,2,3,4\}$ corresponding to the response category (Eisenhauer et al., 2009), or genotype data with covariate taking values in $\{0,1,2\}$ corresponding to the number of copies of a specified allele. NExON-Bayes builds on the standard graphical spike-and-slab model of Wang (2015), and introduces a sub-model that characterises the dependence between edge inclusion probabilities and sample-level ordinal covariate data, rather than modelling the entries of the precision matrix, as done in other covariate-dependent approaches. This modelling framework allows us to develop a deterministic inference algorithm that is more scalable than existing alternatives in the literature.
@@
This paper is organised as follows. Section 2 (Methods) introduces the terminology and notation used throughout the article, our proposed model, and the variational inference algorithm. Next, Section 3 (Results) assesses the performance of our approach by benchmarking the model against several competitors through simulations, and investigates its benefits in a real-life setting where we estimate the proteomic networks of breast carcinoma patients. Finally, Section 4 (Discussion) summarises our contributions and avenues for extensions of this work.
@@ 2.1 Terminology and notation
where $\boldsymbol{\Omega}$ is the positive definite inverse covariance matrix, known as the precision matrix. In this work, we consider settings in which each of the $N$ samples is characterised by a sample-level ordinal covariate $a$ taking values in a set $\mathcal{A}$. To leverage this information, we assume sample homogeneity only within groups defined by equal covariate value, rather than across all samples. Specifically,
@@
where $\boldsymbol{\Omega}^{(a)}$ is the precision matrix specific to the $N_{a}$ samples with covariate value $a\in\mathcal{A}$. [^1] When referring to the covariate value, we use ‘$a$’, and when referring to networks or data that correspond to a covariate value $a$ we use ‘$(a)$’. A covariate-specific dataset can then be denoted as $\boldsymbol{Y}^{(a)}=(\boldsymbol{y}^{(a)}_{1},...,\boldsymbol{y}^{(a)}_{N_{a}})$.
@@ 2.2 Graphical spike-and-slab model
[…]
Figure 1: Schematic representation of the NExON-Bayes model. Shaded nodes are observed, and non-shaded are latent variables that are inferred.
To induce sparsity, our proposed approach, NExON-Bayes, relies on a continuous spike-and-slab prior on the off-diagonal elements of each $\mathbf{\Omega}^{(a)}$, similarly as in wang2015scaling. This is achieved by using a mixture of two Gaussian distributions, one with very low variance (the spike) and the other with high variance (the slab). A binary edge inclusion indicator is also introduced within this prior, which is denoted as $\delta_{ij}^{(a)}$, corresponding to the $ij$’th entry of $\mathbf{\Omega}^{(a)}$. Edge inclusion is indicated by $\delta_{ij}^{(a)}=1$, while edge exclusion is indicated by $\delta_{ij}^{(a)}=0$. The spike-and-slab prior is thus represented as
@@
for $1\leq i<j\leq P$. A symmetric matrix is formed by mirroring this lower triangle across the diagonal. Note that a mixture-of-Gaussians spike-and-slab prior is used, following wang2015scaling, rather than the alternative option of a point-mass spike. This choice preserves positive definiteness in the precision matrix during iterative inference through conjugate updating (see Supplementary Material 1.1 & 1.1.1; george1997approaches).
@@ 2.3 Dependence on ordinal covariates
where $\Phi(\cdot)$ is the standard normal cumulative distribution function and $a$ is the ordinal covariate. The parameter $\beta_{ij}$ is an edge-specific regression coefficient quantifying the influence of the covariate on the inclusion of edge $(i,j)$. Specifically, $\beta_{ij}=0$ means no relationship between covariate value and edge inclusion probability, i.e., edge $(i,j)$ has the same probability of being present or absent in all networks. A positive value of $\beta_{ij}$ means that edge inclusion is more likely in networks with high covariate values and vice versa. In addition, $\zeta_{ij}$ captures the varying baseline inclusion probability of edges and acts as an offset within the probit submodel. We assign a Gaussian prior distribution to $\beta_{ij}$,
@@
Whilst a full multivariate normal could be maintained for all $\beta$ variables, this mean-field approximation reduces the computational complexity of the algorithm, aiding scalability to high-dimensional settings by reducing runtime and memory requirements. Moreover, the model’s hierarchical structure implicitly preserves some dependency via the precision matrices (see Figure 2 and analysis in Section 3.2 (Results)). Each $\zeta_{ij}$ is also assigned a Gaussian prior,
@@
where the parameters $n_{0}$ and $t_{0}^{2}$ are fixed following 10.1093/biostatistics/kxae021 to prior beliefs about the mean and variance of the number of edges in the networks, and thus the overall sparsity. The hierarchical structure of these priors is presented schematically in Figure 1.
@@ 2.4 Variational inference algorithm
Stochastic sampling algorithms like Markov chain Monte Carlo (MCMC) are typically computationally demanding for large graphical models (wang2015scaling). Instead, we employ deterministic inference algorithms which have been successfully applied in graphical modelling contexts, see, e.g., li2019expectationconditionalmaximizationapproach, lingjaerde2024scalable, 10.1093/biostatistics/kxae021. Specifically, we develop a variational Bayes expectation conditional maximisation (VBECM) algorithm. The true posterior distribution can be written as $p(\underline{\boldsymbol{\Omega}},\boldsymbol{\Theta}\mid\boldsymbol{y})$, where $\underline{\boldsymbol{\Omega}}=\{\boldsymbol{\Omega}^{(a)}\}_{a\in\mathcal{A}}$ and $\boldsymbol{\Theta}$ is the set of all variables outside of the precision matrix. This true distribution is approximated by using a variational distribution with mean-field approximation:
$q(\underline{\boldsymbol{\Omega}},\boldsymbol{\Theta})=\prod_{a\in\mathcal{A}}q(\boldsymbol{\Omega}^{(a)})\prod_{a\in\mathcal{A}}\prod_{i<j}q(\delta^{(a)}_{ij},z_{ij}^{(a)})\prod_{i<j}q(\zeta_{ij})q(\sigma^{-2})\prod_{i<j}q(\beta_{ij}),$
@@
$\mathcal{L}(q)=\mathbb{E}_{q(\underline{\boldsymbol{\Omega}},\boldsymbol{\Theta})}\{\log p(\boldsymbol{Y},\underline{\boldsymbol{\Omega}},\boldsymbol{\Theta})\}-\mathbb{E}_{q(\underline{\boldsymbol{\Omega}},\boldsymbol{\Theta})}\{\log q(\underline{\boldsymbol{\Omega}},\boldsymbol{\Theta})\}$
(bishop2006pattern). The optimal distribution for the $n$-th factor of $\boldsymbol{\Theta}$, denoted $q_{n}(\boldsymbol{\Theta}_{n})$ is found by iteratively taking the expectation of the logarithm of the joint distribution over all the other factors,
@@
where the constant absorbs all terms not depending on $\boldsymbol{\Theta}_{n}$. Inferring the precision matrices requires more attention, as the variational distribution of each $\boldsymbol{\Omega}^{(a)}$ is intractable and therefore necessitates a conditional-maximisation step where each $\boldsymbol{\Omega}^{(a)}$ is updated in a blockwise fashion according to wang2015scaling. Importantly, this updating scheme ensures that positive definiteness is maintained at each iteration, as detailed in Supplementary Material 1.1.1. The update rules for each variable, along with their derivations and ELBOs are given in supplementary Material 1.1 and 1.2.
Finally, we use the extended Bayesian Information Criteria ($\text{BIC}_{\gamma}$) to select the spike variance $\nu_{0}$. $\text{BIC}_{\gamma}$ extends the Bayesian Information Criteria (BIC) with an additional regularisation term . The $\text{BIC}_{\gamma}$ of a network-specific graphical estimation, $\boldsymbol{\Omega}^{(a)}_{est.}$, can be found using the following equation:
@@
where $\ell(\cdot)$ is the log likelihood and $|E^{(a)}|$ is the number of edges present in estimated network $(a)$ (foygel2010extended). The regularisation parameter $\gamma$ controls the sparsity of the estimates. chen2008extended have shown that, provided that $0\leq\gamma\leq 1$, the asymptotic properties of $\text{BIC}_{\gamma}$ are maintained and valid model selection is achieved; we follow the default setting employed by these authors, namely $\gamma=0.5$, which balances model complexity and selection accuracy in high-dimensional contexts (foygel2010extended). To optimise the values of $\nu_{0}$, we use line searches to find a value of $\nu_{0}^{(a)}$ that minimises $\text{BIC}_{\gamma}(\boldsymbol{\Omega}_{est.}^{{(a)}})$ for each network $(a)$ in the vanilla graphical spike-and-slab model, which gives a strong estimate for expected sparsity.
@@ 3.1 Simulated Data
We use simulation studies to assess the performance of NExON-Bayes by analyzing how well it can recover the true precision matrix from which simulated data is generated. In the main scenario that we consider, linear relations are simulated between probit edge inclusion probabilities and ordinal covariates taking values $a\in\mathcal{A}=\{1,2,3,4\}$. We correspondingly generate four precision matrices, each with $P=100$ nodes, from which we sample normally distributed data according to Equation (1), where the number of observations for each dataset is $N_{a}=150$. To start, network $(1)$ is simulated as scale-free, with $100$ edges ($|E|=P$) and thus a sparsity of $0.02$. ‘Scale-free’ refers to a graph that has a degree distribution that follows a power law, which leads to few nodes having several connected edges (hubs) and most nodes having very few connected edges. This trait emulates generic biological networks, which are often scale-free (barabasi2003scale). The other networks are then created by altering the entries of this first network. Specifically, the zero-valued partial correlations corresponding to $50$ edges that are not present in network $(1)$ are simulated to linearly increase with the value of the covariate $a$. Similarly, the non-zero valued partial correlations corresponding to $50$ edges that are present in network $(1)$ are simulated to linearly diminish to zero, such that they are not present ($\rho_{ij}=0$) in network $(4)$. The remaining $50$ edges present in network $(1)$ are present in all four networks, with unchanged partial correlations. We call edges with increasing partial correlation ‘appearing’ edges and those with decreasing partial correlation ‘disappearing’ edges.
Simulating precision matrices in this way gives rise to an interesting artefact: appearing and disappearing edges have smaller precision matrix entries in networks $(2)$ and $(3)$ and are thus more difficult to uncover. A more detailed description of how these networks are generated can be found in Supplementary Material 2.1 and seen schematically in Supplementary Figure S1.
@@ 3.2 Borrowing information across networks using joint modelling
Figure 2: Posterior mean estimates of each $\beta_{ij}$ (lower diagonal part) versus simulated coefficients of the ordinal covariate (upper diagonal part). Note that there is no real ‘true’ $\mathbf{\beta}$, as the precision matrices are not simulated from the model. However, there is a known structure based on positive ($\beta_{ij}>0$) and negative ($\beta_{ij}<0$) dependencies between covariate and partial correlation. The outline of the true structure is overlain on the estimated structure to ensure a clear comparison.
We now use this simulated data to compare the performance of NExON-Bayes to the standard single-network Bayesian graphical spike-and-slab model (SSL) (10.1093/biostatistics/kxae021; li2019expectationconditionalmaximizationapproach), upon which our extension is built. Specifically, we evaluate whether introducing a direct relationship between a covariate and the posterior probability of edge inclusion (PPI) indeed boosts statistical power – as a first sanity check that the model behaves as it is meant to.
@@
The level of information sharing is controlled by the edge-specific regression coefficient $\beta_{ij}$. Not only is the estimation of $\beta_{ij}$ useful in increasing statistical power, it also provides insight on sample-level heterogeneity in terms of the effect of the sample-level covariate on the graphical structure. We now consider a smaller simulated setting (with $P=20$) focused on assessing the ability of NExON-Bayes to recover these edge-specific regression coefficients. The results shown in Figure 2 indicate that, while the model does make a mean-field assumption of each $\beta_{ij}$, NExON-Bayes successfully captures the changes of edges associated with covariates, information which, as we will illustrate in the cancer data analysis (see Real-World Application), can enable clinically meaningful interpretation of the estimated network structures. Such insights are not captured intrinsically by state-of-the-art single or joint network modelling approaches.